Savannah has the following incomes for the tax year 2019-20. Gross salary of £12,800 (PAYE deducted £360); building society interest of £3,050 and interest from company debentures of £2,500.
Calculate Savannah’s income tax payable for tax year 2019-20
1
Expert's answer
2020-07-08T17:47:49-0400

2019-2020:


UK income tax rates for 2019–2020

Income Tax Rate

from 0 to 12 500 £ 0%

from 12 501 to 50 000 £ 20%

from 50,001 to 150,000 £ 40%

from 150 001 £ 45%


1)12800-360=12440

"12 440*0=0"


2)building society interest £ 5,000 do not have to pay tax


3) 12440+2500=14940

base rate

2500-2000=500

"500\times0.075=37.5"


0+0+37.5=37.5